OpenStudy (anonymous):

Simplify the number using the imaginary unit i. square root -28

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

\[\sqrt{-28}=\sqrt{-1}\sqrt{28}\] we know that \[i=\sqrt{-1}\]so no we have \[i\sqrt{28}\] can you simplify 28 (hint _ * 6)

OpenStudy (anonymous):

take root of 28 and put sqrt-1=i

OpenStudy (anonymous):

(_ * 7)

OpenStudy (anonymous):

7 * 4 .. 4 is a perfect square